THE MICROWAVE SPECTRUM AND MOLECULAR STRUCTURE OF 2-CHLORO-1,1-DIFLUOROETHYLENE AND PROGRESS TOWARDS ITS COMPLEXES WITH ARGON AND HY-DROGEN FLUORIDE

\underline{MARK D. MARSHALL}, HELEN O. LEUNG, and JASMINA C. CHEUNG-LAU, Department of Chemistry, Amherst College, P.O. Box 5000, Amherst, MA 01002-5000.

Substitution of fluorine for hydrogen in 1,1-difluoroethylene leads to the observation of unexpected, but ultimately understandable, structures for the complexes of 1,1,2-trifluoroethylene with both hydrogen fluoride^{a} and acetylene.^{b} As a first step towards investigating the effect of chlorine substitution on the same carbon atom, it is necessary to refine the molecular structure of 2-chloro-1,1-difluoroethylene available in the literature.^{c} Fourier transform microwave spectra of six isotopomers of 2-chloro-1,1-difluoroethylene are obtained in the 7–21 GHz region. Strong b-type and weaker a-type transitions, which are split by the chlorine nuclear quadrupole hyperfine interaction, are observed. Kraitchman substitution coordinates, fits to the molecular moments of inertia, and ab initio results are used to determine bond lengths and angles for the molecule of sufficient quality to be used in determining the structures of its complexes. Ab initio calculations on the complexes of 2-chloro-1,1-difluoroethylene with argon and with hydrogen fluoride are used to assist in the analysis of spectra obtained in pulsed jet expansions of mixtures of 2-chloro-1,1-difluoroethylene with hydrogen fluoride in both argon and first run neon.

aH. O. Leung and M. D. Marshall, J. Chem. Phys. \underline{126}, in press (2006).

bH. O. Leung, M. D. Marshall, and W. T. Cashion, 61^{st} Ohio State University International Symposium on Molecular Spectroscopy, RE12 (2006).

cD. R. Jenkins, and T. M. Sugden, Trans. Faraday Soc. \underline{55}, 1473 (1959); S. Chandra, J. Phys. Chem. \underline{71}, 1927 (1967).